# Maintaining Distributed Data Structures in Dynamic Peer-to-Peer Networks

### v1

@@ Abstract
We study robust and efficient distributed algorithms for building and maintaining distributed data structures in dynamic Peer-to-Peer (P2P) networks. P2P networks are characterized by a high level of dynamicity with abrupt heavy node churn (nodes that join and leave the network continuously over time). We present a novel algorithm that builds and maintains with high probability a skip list for $poly(n)$ rounds despite $\mathcal{O}(n/\log n)$ churn per round ($n$ is the stable network size). We assume that the churn is controlled by an oblivious adversary (that has complete knowledge and control of what nodes join and leave and at what time and has unlimited computational power, but is oblivious to the random choices made by the algorithm). Moreover, the maintenance overhead is proportional to the churn rate. Furthermore, the algorithm is scalable in that the messages are small (i.e., at most $\text{polylog}{(n)}$ bits) and every node sends and receives at most $\text{polylog}{(n)}$ messages per round.
Our algorithm crucially relies on novel distributed and parallel algorithms to merge two $n$-elements skip lists and delete a large subset of items, both in $\mathcal{O}(\log n)$ rounds with high probability. These procedures may be of independent interest due to their elegance and potential applicability in other contexts in distributed data structures.
To the best of our knowledge, our work provides the first-known fully-distributed data structure that provably works under highly dynamic settings (i.e., high churn rate). Furthermore, they are localized (i.e., do not require any global topological knowledge). Finally, we believe that our framework can be generalized to other distributed and dynamic data structures including graphs, potentially leading to stable distributed computation despite heavy churn.
@@ 1.1 Model: Dynamic Networks with Churn
We consider a synchronous dynamic network controlled by an oblivious adversary, i.e., the adversary does not know the random choices made by the nodes. The adversary fixes a dynamically changing sequence of sets of network nodes $\mathcal{V}=\left(V_{0},V_{1},V_{2},\dots,V_{t},\dots\right)$ where $V_{t}\subset U$, for some universe of nodes $U$ and $t\geq 1$, denotes the set of nodes present in the network during round $t$. A node $u$ such that $u\in V_{t}$ and $u\notin V_{t+1}$ is said to be leaving at time $t+1$. Similarly, a node $v\notin V_{t}$ and $v\in V_{t+1}$ is said to be joining the network at time $t+1$. Each node has a unique ID and we simply use the same notation (say, $u$) to denote both the node as well as its ID. The lifetime of a node $u$ is (adversarially chosen to be) a pair $(s_{u},t_{u})$, where $s_{u}$ refers to its start time and $t_{u}$ refers to its termination time. The size of the vertex set is assumed to be stable $|V_{t}|=n$ for all $t$; this assumption can be relaxed to consider a network that can shrink and grow arbitrarily as discussed in Section 2.1. Each node in $U$ is assumed to have a unique ID chosen from an $ID$ space of size polynomial in $n$. Moreover, for the first $B=\beta\log n$ rounds (for a sufficiently large constant $\beta>0$), called the bootstrap phase, the adversary is silent, i.e., there is no churn, more precisely, $V_{0}=V_{1}=\dots=V_{B}$. We can think of the bootstrap phase as an initial period of stability during which the protocol prepares itself for a harsher form of dynamism. Subsequently, the network is said to be in maintenance phase during which $\mathcal{V}$ can experience churn in the sense that a large number of nodes might join and leave dynamically at each time step.
Communication is via message passing. Nodes can send messages of size $\mathcal{O}(\log n)$ bits to each other if they know their IDs, but no more than $\mathcal{O}(\text{polylog}{(n)})$ incoming and outgoing messages are allowed at each node per round. Furthermore, nodes can create and delete edges over which messages can be sent/received. This facilitates the creation of structured communication networks. A bidirectional edge $e=(u,v)$ is formed when one of the end points $u$ sends $v$ an invitation message followed by an acceptance message from $v$. The edge $e$ can be deleted when either $u$ or $v$ sends a delete message. Of course, $e$ will be deleted if either $u$ or $v$ leaves the network.
@@ 1.2 Problem Statement
Our primary goal is to build and maintain a data structure of all the items/nodes in the network. The data structure takes at most $I$ rounds to insert a data item $u$ and at most $R$ rounds to remove it. $I$ and $R$ are parameters that we would like to minimize. The distributed data structure should be able to answer membership queries. Specifically, each query $q(x,r,s)$ is initiated at some source node $s$ in round $r$ and asks whether data item $x$ is present in the network currently. This implies that node $s$ now wishes to know if some node in the network has the value $x$. The query should be answered by round $r+Q$ where $Q$ is the query time, i.e., the time to respond to queries. If $t_{s}>r+Q$, i.e., node $s$ is in the network until round $r+Q$, it must receive the response. We are required to give the guarantee that the query will be answered correctly as long as either (i) there is a node with associated value $x$ whose effective lifetime subsumes the time range $[r,r+Q]$ (in which case the query must be answered affirmatively) or (ii) there is no node with value $x$ whose effective lifetime has any overlap with $[r,r+Q]$ (in which case the query must be answered negatively). In all other cases, we allow queries to be answered incorrectly.
In addition, all our algorithms must satisfy a dynamic notion of resource-competitiveness [19]. Let $T$ be any interval between two time instants $t_{s}$ and $t_{e}$. We require the work in the interval $T$ to be proportional (within $\text{polylog}{(n)}$ factors) to the amount of churn experienced from time $t_{s}-\mathcal{O}(\log n)$ to $t_{e}$. Formally, define the amount of work $W_{t}$ as the overall number of exchanged messages plus newly formed edges among nodes at time $t$. Let $\mathcal{W}(t_{s},t_{e})=\sum_{i={t_{s}}}^{t_{e}}W_{i}$ and $C(t_{s},t_{e})$ to be the amount of work and churn experienced in the interval of time between $t_{s}$ and $t_{e}$, respectively. We require that $\mathcal{W}(t_{s},t_{e})\in\tilde{\mathcal{O}}\left(C(t_{s}-\mathcal{O}(\log n),t_{e})\right)$ w.h.p. for any interval of time $t_{s}$, $t_{e}$, where with the notation $\tilde{\mathcal{O}}(\cdot)$ we ignore $\text{polylog}{(n)}$ factors. Formally,
@@
In this work, whenever we refer to an algorithm as dynamic resource competitive, we mean that the algorithm is $(\alpha,\beta)$-dynamic resource competitive with $\alpha=\mathcal{O}(\log n)$ and $\beta=\text{polylog}{(n)}$.
@@ 1.3 Our Contributions
Any query to a data structure will require some response time, which we naturally wish to minimize. Furthermore, queries will be imprecise to some extent if the data structure is dynamic. To see why, consider a membership query of the form “is node $u$ in the dynamic skip list?” initiated in some round $r$ asking if a node $u$ is in the network or not. Consider a situation where $u$ is far away from the node at which the query was initiated. Suppose $u$ is then churned out shortly thereafter. Although $u$ was present at round $r$, it may or may not be gone when the query procedure reached $u$. Such ambivalences are inevitable, but we wish to limit them. Thus, we define an efficiency parameter $Q$ such that (i) queries raised at round $r$ are answered by round $r+Q$ and (ii) response must be correct in the sense that it must be “Yes” (resp., “No”) if $u$ was present (resp., not present) from round $r$ to $r+Q$. If $u$ was only present for a portion of the time between $r$ and $r+Q$, then either of the two answers is acceptable. Quite naturally, we wish to minimize $Q$ and for our dynamic data structure, we show that $Q\in\mathcal{O}(\log n)$.
@@
    A novel algorithm that constructs and maintains a skip list in a dynamic P2P network with an adversarial churn rate up to $\mathcal{O}(n/\log n)$ per round.
@@
To the best of our knowledge, our approach is novel and fully-distributed skip list data structure that works under highly dynamic settings (high churn rates per step). Furthermore, all the proposed algorithms are localized, easy to implement and scalable. Our major contribution can be summarized in the following theorem.
@@
Given a dynamic set of peers initially connected in some suitable manner (e.g., as a single path) that is stable for an initial period of $\mathcal{O}(\log n)$ rounds (i.e., the so-called bootstrap phase) and subsequently experiencing heavy adversarial churn at a churn rate of up to $\mathcal{O}(n/\log n)$ nodes joining/leaving per round, we
  • provide an $\mathcal{O}(\log n)$ round algorithm to construct a resilient skip list that can withstand heavy adversarial churn at a churn rate of up to $\mathcal{O}(n/\log n)$ nodes joining/leaving per round,
  • describe a fully distributed algorithm that maintains the resilient skip list with every new node inserted into the data structure in $\mathcal{O}(\log n)$ rounds such that membership queries can be answered with efficiency parameter $Q\in\mathcal{O}(\log n)$.
All nodes send and receive at most $\mathcal{O}(\text{polylog}{(n)})$ messages per round, each comprising at most $\mathcal{O}(\text{polylog}{(n)})$ bits. Moreover, our algorithms are dynamically resource competitive according to Definition 1.1. The maintenance protocol ensures that the resilient skip list is maintained effectively for at least $poly(n)$ rounds with high probability.
Implications.
A consequence of our maintenance algorithm is that it opens up opportunities for more general distributed computation in the DNC model. In fact, this paper formally shows how to build and maintain a non-trivial data structure under such a high dynamic settings. We now informally discuss several ideas that illustrate how our framework can be generalized. Of course, significant followup work is required to formally prove our claims. We believe this can be extended to maintaining any pointer-based data structure. In particular, we believe we can maintain an arbitrary graph over the DNC and solve a wide range of fundamental distributed graph computation tasks in the DNC model that, until now, seemed implausible.
Our current approach assumes data items are embedded into the nodes. It might be more feasible to consider the overlay network and the graph structure as two separate entities. This will allow us to retain the graph structure even when nodes in the overlay network are deleted by the adversary.
Another implication is the construction and maintenance of other more sophisticated data structures like skip graphs and its ilk [3, 39, 47]. All these data structures are not resilient to churns, and their maintenance protocols are not fast enough to recover the data structure after the failure of some nodes. Indeed, these protocols (see [3, 39, 47]) might need $\mathcal{O}(n)$ rounds to rebuild the skip-graph and they strictly require no additional churn to be able to fix the data structure. Our maintenance protocol overcomes these problems and provides a $\mathcal{O}(\log n)$ rounds skip graph constructing protocol and a technique able to repair them in $\mathcal{O}(\log n)$ in the presence of an almost linear churn of $\mathcal{O}(n/\log n)$ at every round. Moreover, our maintenance mechanism allows for the users to query the data structure while being maintained.
Finally, some ideas from our results could be used in the centralized batch parallel setting to quickly insert batches of new elements in skip lists (or skip graphs) data structures (see e.g.,[83]) and in the fully dynamic graph algorithms settings (see for example the survey [43]) to perform fast updates of fully dynamic data structures.
High-level Overview and Technical Contributions.
@@
Before delving into the protocol’s description, we highlight one of the key ideas of our paper. As previously mentioned, the network comprises two networks – the overlay and the distributed data structure. However, instead of having only one network for the current skip list data structure we keep three: the first one is a live network on which queries are executed, the second one is a clean version of the live network on which we perform updates after every churn and the latter, is an additional temporary buffer network on which we store the newly added elements. Such a buffer will be promptly merged to the clean network during each cycle. Such a three-network architecture allows us to describe the maintenance process in a clear and simple way. (Note that the nodes are not actually replicated into multiple copies for each of the networks. The same set of nodes will maintain all the networks logically.)
Another key aspect of our protocol is that when the oblivious adversary removes some nodes from the overlay network (thus from the data structures as well), we have that (1) the overlay network maintenance protocol in [13] ensures that it will remain connected with high probability at every round and (2) all the elements that disappeared from the data structures will be temporarily replaced by some surviving group of nodes in the overlay network. Moreover, when (2) happens, all these replacement nodes, in addition to answering their own queries, they also answer the ones addressed to the nodes they are covering for. This is possible since we can always assume some level of redundancy in the overlay network. For example, we can assume that a node $u$ in the overlay network has an updated copy of all its neighbors’ values and pointers in the data structure. (Recall that both network are designed to be low-degree networks).
We now describe a cycle of our maintenance protocol which consists of four major phases. Moreover, we can assume that at the beginning of the cycle, live and clean networks contain the same elements. Next, assume that the adversary replaced $\mathcal{O}(n/\log n)$ from the network, this means that $\mathcal{O}(n/\log n)$ nodes have been removed from the networks (overlay, live and clean) and new $\mathcal{O}(n/\log n)$ have been added to the overlay network. After such a churn, there are some (old) nodes in the overlay network covering for the removed ones in both clean and live networks.
In the first phase, we quickly remove from the clean network those nodes that have been covered by other surviving nodes in the overlay network. In Section 2.3, we provide an $\mathcal{O}(\log n)$ rounds protocol that successfully polish the clean network.
@@
The above maintenance cycle maintains the distributed data structure with probability at least $1-\frac{1}{n^{c}}$ for some arbitrarily big constant $c\geq 1$. This implies that the expected number of cycles we have to wait before getting the first failure is $n^{c}$ and that the probability that our protocol correctly maintains the data structure for some $r<c$ rounds is at least $1-\frac{1}{n}$.
Finally, we show how to extend our maintenance cycle to skip graphs-like data structures (Section 2.7) and that our technique can be easily generalized to handle cases in which each node in the overlay network possesses more than one element in the data structure (Section 2.8).
1.4 Related Works

### v2

# Highly Dynamic and Fully Distributed Data Structures

@@ Abstract
We study robust and efficient distributed algorithms for building and maintaining distributed data structures in dynamic Peer-to-Peer (P2P) networks. P2P networks are characterized by a high level of dynamicity with abrupt heavy node churn (nodes that join and leave the network continuously over time). We present a novel algorithmic framework to build and maintain, with high probability, a skip list for $\text{poly}(n)$ rounds despite a churn rate of $O(n/\log n)$, which is the number of nodes joining and/or leaving per round; $n$ is the stable network size. We assume that the churn is controlled by an oblivious adversary that has complete knowledge and control of what nodes join and leave and at what time and has unlimited computational power, but is oblivious to the random choices made by the algorithm. Importantly, the maintenance overhead in any interval of time (measured in terms of the total number of messages exchanged and the number of edges formed/deleted) is (up to log factors) proportional to the churn rate. Furthermore, the algorithm is scalable in that the messages are small (i.e., at most $\text{polylog}(n)$ bits) and every node sends and receives at most $\text{polylog}(n)$ messages per round.
To the best of our knowledge, our work provides the first-known fully-distributed data structure and associated algorithms that provably work under highly dynamic settings (i.e., high churn rate that is near-linear in $n$). Furthermore, the nodes operate in a localized manner.
Our framework crucially relies on new distributed and parallel algorithms to merge two $n$-element skip lists and delete a large subset of items, both in $\mathcal{O}(\log n)$ rounds with high probability. These procedures may be of independent interest due to their elegance and potential applicability in other contexts in distributed data structures.
Finally, we believe that our framework can be generalized to other distributed and dynamic data structures including graphs, potentially leading to stable distributed computation despite heavy churn.
@@ 1.1 Model: Dynamic Networks with Churn
We consider a synchronous dynamic network controlled by an oblivious adversary, i.e., the adversary does not know the random choices made by the nodes. The adversary fixes a dynamically changing sequence of sets of network nodes $\mathcal{V}=\left(V_{0},V_{1},V_{2},\dots,V_{t},\dots\right)$ where $V_{t}\subset U$, for some universe of nodes $U$ and $t\geq 1$, denotes the set of nodes present in the network during round $t$. A node $u$ such that $u\in V_{t}$ and $u\notin V_{t+1}$ is said to be leaving at time $t+1$. Similarly, a node $v\notin V_{t}$ and $v\in V_{t+1}$ is said to be joining the network at time $t+1$. Each node has a unique ID and we simply use the same notation (say, $u$) to denote both the node as well as its ID. The lifetime of a node $u$ is (adversarially chosen to be) a pair $(s_{u},t_{u})$, where $s_{u}$ refers to its start time and $t_{u}$ refers to its termination time. The size of the vertex set is assumed to be stable $|V_{t}|=n$ for all $t$; this assumption can be relaxed to consider a network that can shrink and grow arbitrarily as discussed in Section 2.1. Each node in $U$ is assumed to have a unique ID chosen from an $ID$ space of size polynomial in $n$. Moreover, for the first $B=\beta\log n$ rounds (for a sufficiently large constant $\beta>0$), called the bootstrap phase, the adversary is silent, i.e., there is no churn; more precisely, $V_{0}=V_{1}=\dots=V_{B}$. We can think of the bootstrap phase as an initial period of stability during which the protocol prepares itself for a harsher form of dynamism. Subsequently, the network is said to be in maintenance phase during which $\mathcal{V}$ can experience churn in the sense that a large number of nodes might join and leave dynamically at each time step.
Communication is via message passing. Nodes can send messages of size $\mathcal{O}(\text{polylog}{(n)})$ bits to each other if they know their IDs, but no more than $\mathcal{O}(\text{polylog}{(n)})$ incoming and outgoing messages are allowed at each node per round. Furthermore, nodes can create and delete edges over which messages can be sent/received. This facilitates the creation of structured communication networks. A bidirectional edge $e=(u,v)$ is formed when one of the end points $u$ sends $v$ an invitation message followed by an acceptance message from $v$. The edge $e$ can be deleted when either $u$ or $v$ sends a delete message. Of course, $e$ will be deleted if either $u$ or $v$ leaves the network.
@@ 1.2 Problem Statement
Our primary goal is to build and maintain an approximate distributed data structure of all the items/nodes in the network. The data structure takes at most $I$ rounds to insert a data item $u$ and at most $R$ rounds to remove it. $I$ and $R$ are parameters that we would like to minimize. The distributed data structure should be able to answer membership queries. Specifically, each query $q(x,r,s)$ is initiated at some source node $s$ in round $r$ and asks whether data item $x$ is present in the network currently. This implies that node $s$ now wishes to know if some node in the network has the value $x$. The query should be answered by round $r+Q$ where $Q$ is the query time, i.e., the time to respond to queries. If $t_{s}>r+Q$, i.e., node $s$ is in the network until round $r+Q$, it must receive the response. We are required to give the guarantee that the query will be answered correctly as long as either (i) there is a node with associated value $x$ whose effective lifetime subsumes the time range $[r,r+Q]$ (in which case the query must be answered affirmatively) or (ii) there is no node with value $x$ whose effective lifetime has any overlap with $[r,r+Q]$ (in which case the query must be answered negatively). In all other cases, we allow queries to be answered arbitrarily.
Unlike classical data structures, which guarantee strong consistency, our construction deliberately prioritizes scalability and liveness under high churn. As a result, it provides eventual consistency, meaning that queries will reflect the correct state once the system has stabilized, but temporary inconsistencies may arise during periods of high churn. We refer to this design as an approximate distributed data structure, reflecting its structural inspiration and relaxed consistency guarantees. This tradeoff is necessary to maintain efficiency and fault-tolerance in the presence of high adversarial churn.
[…]
Figure 1: Visual representation of our definition of dynamic work efficiency.
@@
Figure 1 is a visual representation of the $(\alpha,\beta)$-dynamic resource competitiveness. The red area represents the amount of churn experienced over time while the green area represent the workload of the algorithm to cope with the churn experienced so far. Figure 1 illustrates a spike in churn rate just prior to $t_{s}$ that resulted in an increased workload after $t_{s}$. Thus, to capture such delayed effects, the dynamic resource competitiveness definition considers $C(t_{s}-\alpha,t_{e}))$ instead of just $C(t_{s},t_{e}))$. In this work, whenever we refer to an algorithm as dynamic resource competitive, we mean that the algorithm is $(\alpha,\beta)$-dynamic resource competitive with $\alpha=\mathcal{O}(\log n)$ and $\beta=\text{polylog}{(n)}$.
@@ 1.3.1 Main Result
Any query to a data structure will require some response time, which we naturally wish to minimize. Furthermore, queries will be imprecise to some extent if the data structure is dynamic. To see why, consider a membership query of the form “is node $u$ in the dynamic skip list?” initiated in some round $r$ asking if a node $u$ is in the network or not. Consider a situation where $u$ is far away from the node at which the query was initiated. Suppose $u$ is then deleted shortly thereafter. Although $u$ was present at round $r$, it may or may not be gone when the query procedure reached $u$. Such ambivalences are inevitable, but we wish to limit them. Thus, we define an efficiency parameter $Q$ such that (i) queries raised at round $r$ are answered by round $r+Q$ and (ii) response must be correct in the sense that it must be “Yes” (resp., “No”) if $u$ was present (resp., not present) from round $r$ to $r+Q$. If $u$ was only present for a portion of the time between $r$ and $r+Q$, then either of the two answers is acceptable. Quite naturally, we wish to minimize $Q$ and for our dynamic data structure, we show that $Q\in\mathcal{O}(\log n)$.
@@
    An algorithm that constructs and maintains a skip list in a dynamic P2P network with an adversarial churn rate up to $\mathcal{O}(n/\log n)$ per round.
@@
To the best of our knowledge, our approach is the first fully-distributed skip list data structure that works under highly dynamic settings (high churn rates per step). Furthermore, all the proposed algorithms are localized, easy to implement and scalable. Our major contribution can be summarized in the following theorem.
@@
Given a dynamic set of peers initially connected in some suitable manner (e.g., as a single path) that is stable for an initial period of $\mathcal{O}(\log n)$ rounds (i.e., the so-called bootstrap phase) and subsequently experiencing heavy adversarial churn at a churn rate of up to $\mathcal{O}(n/\log n)$ nodes joining/leaving per round, there exists a resilient skip list, a distributed data structure that can withstand heavy adversarial churn at a rate of up to $\mathcal{O}(n/\log n)$ nodes joining/leaving per round. We provide the following algorithms to support this data structure.
  • An $\mathcal{O}(\log n)$ round algorithm to construct the resilient skip list during the bootstrap phase,
  • a fully distributed algorithm that maintains the resilient skip list as nodes are inserted into or deleted from the data structure with the guarantee that the data structure reflects the updates within $\mathcal{O}(\log n)$ rounds,
  • and a membership query algorithm that any peer can invoke and can be answered with efficiency parameter $Q\in\mathcal{O}(\log n)$.
All nodes send and receive at most $\mathcal{O}(\text{polylog}{(n)})$ messages per round, each comprising at most $\mathcal{O}(\text{polylog}{(n)})$ bits. Moreover, such a maintenance algorithm is dynamically resource competitive according to Definition 1.1. The maintenance protocol ensures that the resilient skip list is maintained effectively for at least $n^{d}$ rounds with high probability, where $d\geq 1$ can be an arbitrarily chosen constant.
1.3.2 High-level Overview and Technical Contributions
@@
Before delving into the protocol’s description, we highlight one of the ideas of our paper. As previously mentioned, the network comprises two networks – the overlay and the distributed data structure. However, instead of having only one network for the current skip list data structure we keep three: the first one is a live network on which queries are executed, the second one is a clean version of the live network on which we perform updates after joins and leaves and the latter, is an additional temporary buffer network on which we store the newly added elements. Such a buffer will be promptly merged to the clean network during each cycle. This three-network architecture allows us to describe the maintenance process in a clear and simple way. (Note that the nodes are not actually replicated into multiple copies for each of the networks. The same set of nodes will maintain all the networks logically.)
Another key aspect of our protocol is that when the oblivious adversary removes some nodes from the overlay network (thus from the data structures as well), we have that (1) the overlay network maintenance protocol in [13] ensures that it will remain connected with high probability at every round and (2) all the elements that disappeared from the data structures will be temporarily replaced by some surviving group of nodes in the overlay network. Moreover, when (2) happens, all these replacement nodes, in addition to answering their own queries, they also answer the ones addressed to the nodes they are covering for. This is possible since we can always assume some level of redundancy in the overlay network. For example, we can assume that a node $u$ in the overlay network has an updated copy of all its neighbors’ values and pointers in the data structure. (Recall that both network are designed to be low-degree networks.)
We now describe a cycle of our maintenance protocol which consists of four major phases. Moreover, we can assume that at the beginning of the cycle, live and clean networks contain the same elements. Next, assume that the adversary replaced $\mathcal{O}(n/\log n)$ nodes from the network, this means that $\mathcal{O}(n/\log n)$ nodes have been removed from the networks (overlay, live and clean) and new $\mathcal{O}(n/\log n)$ have been added to the overlay network. After the churn, there are some (old) nodes in the overlay network covering for the removed ones in both clean and live networks.
Delete.
In the first phase, we promptly remove from the clean network all nodes that have been covered by surviving nodes in the overlay. In Section 2.3, we describe an $\mathcal{O}(\log n)$-round protocol that efficiently refines and updates the clean network.
Buffer Creation.
@@
The above maintenance cycle maintains the distributed data structure with probability at least $1-\frac{1}{n^{d}}$ for some arbitrarily big constant $d\geq 1$. This implies that the expected number of cycles we have to wait before getting the first failure is $n^{d}$ and that the probability that our protocol correctly maintains the data structure for some $r<d$ rounds is at least $1-\frac{1}{n}$.
Beyond skip lists, we show how our techniques generalize to more sophisticated data structures such as skip graphs (Section 2.7), and how they can be extended to settings where each node stores multiple elements in the data structure (Section 2.8).
Towards a General Framework.
Our contributions naturally lead to a broader definition: in Section 3.2, we introduce a general framework for distributed data structures under churn and define the class of $t$-maintainable data structures, that is, distributed data structures that can be maintained by a continuous maintenance cycle in $\mathcal{O}(t)$ rounds (per cycle) with high probability despite a churn rate of $\mathcal{O}(n/t)$ per round. In this terminology, our skip list maintenance protocol demonstrates that skip lists and skip graphs are $\mathcal{O}(\log n)$-maintainable in the DNC model under near-linear adversarial churn. We believe this framework lays the foundation for a new complexity class of distributed data structures resilient to adversarial churn, enabling future exploration of which classical structures (e.g., trees, graphs) admit efficient $t$-maintenance algorithms, and what inherent trade-offs exist between maintenance time, redundancy, and robustness.